OpenStudy (anonymous):

A prism is made of cubes which 1/3 of a centimeter on one side. What is the volume in cubic cm?

OpenStudy (anonymous):

The prism is made of 8 cubes in a square

OpenStudy (anonymous):

In a cube all sides are equal length. So in this case all sides of the cube are 1/3cm. To find the volume of one cube (V) you have to times the width by the length by the height. So V = (1/3)*(1/3)*(1/3) = 1/27 cubic centimetres. If the prism is made of 8 cubes than the prisms volume would equal 8 times 1/27 = 8/27 cubic centimetres.